A hammer head, with a mass of 10.0 kg, strikes a nail at a horizontal velocity of 12.2 m/s and comes to rest in a time interval of 0.004 s. Find the impulse imparted to the nail and the average force the hammer exerted on the nail?


1
Expert's answer
2022-04-07T13:10:18-0400

(a)

"\Delta p=m\Delta v\\
=10.0*12.2=122\:\rm kg*m/s"

(b)

"F=\frac{122}{0.004}=30500\:\rm N"